OpenStudy (anonymous):

HELP Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) 2x − 6

OpenStudy (anonymous):

they are equal if u use distributive law on left hand side 2(x-3) = 2*x - 2*3 = 2x - 6 = right hand side